e) lim_((x y) to (0 0)) (sin x)/(y) exists.


1
Expert's answer
2021-03-25T01:57:31-0400

Solution:

"\lim _{\left(x,\:y\right)\to \left(0,\:0\right)}\left(\frac{\sin \left(x\right)}{y}\right)"

We calculate limit along two different paths.

Along "x=0:"

"\lim _{\left(x,\:y\right)\to \left(0,\:0\right)}\left(\frac{\sin \left(0\right)}{y}\right)""=0"

Along "x=y:"

"\lim _{\left(y,\:y\right)\to \left(0,\:0\right)}\left(\frac{\sin \left(y\right)}{y}\right)""=1"

Since, they are unequal, limit diverges or does not exist.